# Machine Learning with Multitype Protected Attributes: Intersectional Fairness through Regularisation

### v1

@@ D.1 Justification of applying gradient descent
by adjusting the model parameters $\Theta$, where $\mathcal{L}$ is a task-specific loss function that is assumed to be differentiable. To enable gradient-based optimisation using AdaHessian, our implementation requires the availability of gradient and Hessian information with respect to $\Theta$ for the full objective. Given that $\mathcal{L}$ is differentiable, we show that our regulariser $\psi$ also satisfies these properties, inspired by Schmidt et al. (2007) and Shalev-Shwartz and Ben-David (2014). This allows efficient second-order optimisation of the regularised loss by leveraging curvature information to accelerate convergence.
@@
At the non-differentiable point where $g_{kl}(\Theta)=0$, we rely on the concept of subgradients:
@@
For second-order information, we use AdaHessian to approximate the diagonal of the Hessian via Hutchinson’s method (Yao et al., 2021). Our objective is piecewise differentiable. Away from points where $g_{kl}(\Theta)\neq 0$, autograd provides exact derivatives and AdaHessian computes Hessian vector products as usual. At $g_{kl}(\Theta)=0$ the true second derivative is not defined; PyTorch returns a particular subgradient and an effective zero curvature at that point. Therefore, the curvature used by AdaHessian at the non-smooth locations should be viewed as a practical approximation rather than a formal second-order quantity.
@@ E.1 Details on hypothesis tests for independence
    Distance correlation $\chi^{2}$-test: This test was developed by Shen et al. (2022). It is based on the distance correlation, which is a standardised version of the distance covariance. Let $\mathbf{x}$ and $\mathbf{y}$ be the corresponding sample of $\mathbf{X}$ and $\mathbf{Y}$ with size $n$. The distance correlation is defined as:
@@
    The second type of permutation test looks at the mutual independence between the model prediction and each protected attribute, as defined in Equation (4.3). Similar to the first test, we compute the $p$-value using the same formula in Equation (E.3), but here the test statistic is based on joint distance covariance: $T_{0}=\widetilde{JdCov}^{2}(\hat{\mathbf{y}},\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ and $T^{(r)}=\widetilde{JdCov}^{2}(\hat{\mathbf{y}},\mathbf{s}_{1}^{(r)},\ldots,\mathbf{s}_{d}^{(r)})$.
@@ Appendix I Further discussions in limited sample size
$\begin{aligned} &\widetilde{dCov}^{2}(Z,(X,Y))\\ &\qquad=\widetilde{dCov}^{2}(Z,X)+\widetilde{dCov}^{2}(Z,Y)+\eta(Z,(X,Y))\\ &\qquad=0.0003+(-0.0005)+0.0472=0.0470.\end{aligned}$ (A.2)
@@
$\begin{aligned} &\widetilde{dCov}^{2}(Z,(X,Y))\\ &\qquad=\widetilde{dCov}^{2}(Z,X)+\widetilde{dCov}^{2}(Z,Y)+\eta(Z,(X,Y))\\ &\qquad=(-0.0006)+(-0.0003)+0.0002=-0.0007.\end{aligned}$ (A.4)

### v2

@@ D.1 Justification of applying gradient descent
by adjusting the model parameters $\Theta$, where $\mathcal{L}(\cdot,\cdot)$ is a task-specific loss function that is assumed to be differentiable. To enable gradient-based optimisation using AdaHessian, our implementation requires the availability of gradient and Hessian information with respect to $\Theta$ for the full objective. Given that $\mathcal{L}$ is differentiable, we show that our regulariser $\psi$ also satisfies these properties, inspired by Schmidt et al. (2007) and Shalev-Shwartz and Ben-David (2014). This allows efficient second-order optimisation of the regularised loss by leveraging curvature information to accelerate convergence.
@@
At the non-differentiable point where $g_{kl}(\Theta)=0$, we rely on the concept of subgradients (Clarke, 1990):
@@
For second-order information, we use AdaHessian to approximate the diagonal of the Hessian via Hutchinson’s method (Yao et al., 2021). Our objective is piecewise differentiable. Away from points where $g_{kl}(\Theta)\neq 0$, autograd provides exact derivatives and AdaHessian computes Hessian vector products as usual. At $g_{kl}(\Theta)=0$ the true second-order derivative is not defined; PyTorch returns a particular subgradient and an effective zero curvature at that point. Therefore, the curvature used by AdaHessian at the non-smooth locations should be viewed as a practical approximation rather than a formal second-order quantity.
@@ E.1 Details on hypothesis tests for independence
    Distance correlation $\chi^{2}$-test: This test was developed by Shen et al. (2022). It is based on the distance correlation, which is a standardised version of the distance covariance as discussed in Appendix A. Let $(\mathbf{x},\mathbf{y})$ be the corresponding sample of size $n$ of $(\mathbf{X},\mathbf{Y})$. The distance correlation is defined as:
@@
    Intuitively, permuting the sample breaks any dependency structure between $\hat{\mathbf{y}}$ and $(\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$. In that case, $T^{(r)}$ should be close to zero. If $p$ is large, which means that many $T^{(r)}$, the distance covariances for supposedly independent predictions and protected attributes, are larger than the distance covariance of the original sample $T_{0}$, then it is very likely that our original sample is already independent.
    The second type of permutation test looks at the mutual independence between the model prediction and each protected attribute, as defined in Equation (4.3). Similar to the first test, we compute the $p$-value using the same formula in Equation (E.3), but here the test statistic is based on joint distance covariance: $T_{0}=\widetilde{JdCov}^{2}(\hat{\mathbf{y}},\mathbf{s}_{1},\ldots,\mathbf{s}_{d})$ and $T^{(r)}=\widetilde{JdCov}^{2}(\hat{\mathbf{y}},\mathbf{s}_{1}^{(r)},\ldots,\mathbf{s}_{d}^{(r)})$, as defined in Equation (B.7).
@@ Appendix I Further discussions in limited sample size
$\begin{aligned} &\widetilde{dCov}^{2}(Z,(X,Y))\\ &\hskip 18.49988pt=\widetilde{dCov}^{2}(Z,X)+\widetilde{dCov}^{2}(Z,Y)+\eta(Z,(X,Y))\\ &\hskip 18.49988pt=0.0003+(-0.0005)+0.0472=0.0470.\end{aligned}$ (A.2)
@@
$\begin{aligned} &\widetilde{dCov}^{2}(Z,(X,Y))\\ &\hskip 18.49988pt=\widetilde{dCov}^{2}(Z,X)+\widetilde{dCov}^{2}(Z,Y)+\eta(Z,(X,Y))\\ &\hskip 18.49988pt=(-0.0006)+(-0.0003)+0.0002=-0.0007.\end{aligned}$ (A.4)